OpenStudy (anonymous):

find the domain of the fuction and express the answer in interval notation: f(x)=4x^2 -7x+3

OpenStudy (anonymous):

The domain is all real numbers

OpenStudy (anonymous):

thas is \[(-\infty,\infty)\]

OpenStudy (anonymous):

how would you show the work for that? like when you are trying to solve for the answer.

OpenStudy (anonymous):

There is no invalid inputs.

OpenStudy (anonymous):

It's just like "show the work of 1+1=2"

OpenStudy (anonymous):

well, its like this, there are no constraints on this polynomial or function

OpenStudy (anonymous):

so its still going to be like solving for x

OpenStudy (anonymous):

in other words you can plug and number in the function and get a valid answer of y value

OpenStudy (anonymous):

thats why the domain is all Real numbers

OpenStudy (anonymous):

so there really isnt a way to show work for it because any real number can work fine

OpenStudy (anonymous):

The way I would show work is to factor this into: (4x-3)(x-1) and state that x = 0 at x = 3/4 and x = 1. These are your terminal points. (Where x = 0), aka x-intercepts. If this were an equation where we had a fraction, say (2x + 1)/(x^2 - 4) the domain would be (-infinity, -2) and (-2, 2) and (2, infinity) because division by zero is undefined and the () brackets denote that the number is not in the set. [ ] denotes the number is in the set. So every number up to + or - 2 in this case would be in the set but not + or - 2. Hope this helps clarify. For free math videos go to http://www.tutorsean.net